OpenStudy (anonymous):

the sum of four consecutive integers is 410

OpenStudy (anonymous):

and what is the problem???

OpenStudy (anonymous):

(x-1)+x+(x+1)+(x+2)=410 where (x-1) is the first number.

OpenStudy (anonymous):

(x-1)+x+(x+1)+(x+2)=410 x-1+x+x+1+x+2=410 4x+2=410 4x=408 x=102 and x-1 , the first number is 101, so the numbers are: 101, 102, 103, 104.